Leo is 5 years older than Pat. In 10 years, Leo will be twice Pat's present age. How old is leo now?
Answer by ikleyn(52800)

.
Using obvious notations, we have these two equations


L = P + 5      (1)    ("Leo is 5 years older than Pat.")

L + 10 = 2P    (2)    ("In 10 years, Leo will be twice Pat's present age.")



From equation (1), express  P = L - 5.  In equation (2),  replace P  by  L-5,  based on it.  You will get


L + 10 = 2*(L-5).


Solve for L :


L + 10 = 2L - 10

10 + 10 = 2L - L

====================>  20 = L.


ANSWER.  Leo is 20 years old now.

Solved.